The area of a rectangle is 240 sq in. The length is twice the width. What is the area of the rectangle?
Found 2 solutions by nerdybill, rwm:
Answer by nerdybill(7384)

The area of a rectangle is 240 sq in. The length is twice the width. What is the area of the rectangle?
.
The problem already gives you the area as 240 sq in.
.
Unless you were looking for the dimensions:
Let w = width
then
2w = length
.
w(2w) = 240
2w^2 = 240
w^2 = 120
w = 10.95 inches (width)
.
length:
2w = 2(10.95) = 21.9 inches

Answer by rwm(914)

L*w=240
L=2w
2w*w=240
w*w=120
w=sqrt(120)
w=sqrt(12*10)
w=2 sqrt(30)=10.95
L=4 sqrt(30)=2*10.95=21.90